21/20x + 1/20x = 9x + 1/10 + 19/20x
I got an answer of - 88 and 1/2. Is this correct? If not, please show me what is.

Answer by stanbon(48546)

(21/20)x + (1/20)x = 9x + 1/10 + (19/20)x
---------------------------------------------------
Multiply thru by 20 to get rid of the fractions:
---
21x + x = 180x + 2 + 19x
---
199x - 22x = -2
177x = -2
x = -2/177
============================
Cheers,
Stan H.